OpenStudy (anonymous):

What is the value of the missing exponent? (Type only the numerical value of the missing exponent. Do not skip any spaces.) (2x3)?=4x6

OpenStudy (anonymous):

6?

OpenStudy (kamille):

oh, firstly, you need to multiply 4x6. What is the answer?

OpenStudy (kamille):

And no, it is not 6

OpenStudy (anonymous):

24

OpenStudy (phi):

is that \[ (2x^3)^? = 4x^6 \]

OpenStudy (anonymous):

its 4x^6

OpenStudy (anonymous):

(2x^3)?=4x^6

OpenStudy (phi):

what is \[ 2x^3 \cdot 2x^3 = ?\]

OpenStudy (anonymous):

4

OpenStudy (phi):

2^2 = 2*2= 4 x^3 * x^3 is the same as x*x*x * x*x*x = x*x*x*x*x*x or in short hand (who want to type that out?) x^6

OpenStudy (anonymous):

so now what

OpenStudy (phi):

do you agree 2x^3 * 2x^3 = 4x^6 how do you rewrite the left side using powers? if it were just A*A= B you know that is A^2= B or (stuff)*(stuff) = (stuff)^2

OpenStudy (anonymous):

(2x^3)?

OpenStudy (phi):

to what power?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

wait no

OpenStudy (phi):

(2x^3)^3 = 2x^3 * 2x^3 * 2x^3 that is not what we want

OpenStudy (anonymous):

2?

OpenStudy (phi):

(2x^3)^2 = 2x^3 * 2x^3

OpenStudy (phi):

and 2x^3 * 2x^3 is 2*2*x^3 * x^3 = 4x^6

OpenStudy (anonymous):

okay so it is 2 correct

OpenStudy (anonymous):

(2x3)^2=4x^6

OpenStudy (phi):

you can multiply it out and check it.

OpenStudy (anonymous):

ok

OpenStudy (phi):

but I did up above.

OpenStudy (anonymous):

What is the standard form of the number written in scientific notation? (DO NOT TYPE COMMAS in your answer.) 3.76* 10^4

OpenStudy (anonymous):

37600

OpenStudy (anonymous):

Am i correct ?

OpenStudy (phi):

yes. you can (sort of) check by changing 37600 back to scientific notation

OpenStudy (anonymous):

okay im right :)